Graph the function using a table of at least 5 values: h(x) = (3x - 4 X
Answer #1

Solution: Given h(x) = e - 4 h(x) s. h(x) = er - 4 4. 0 -3 3 0.1 2. -2.6501412 -2.1778812 0.2 1 10.3 -1.5403969 0.4 -0.679853